# Versatile Multimodal Controls for Whole-Body Talking Human Animation

### v2

@@
Abstract
Human animation from a single reference image shall be flexible to synthesize whole-body motion for either a headshot or whole-body portrait, where the motions are readily controlled by audio signal and text prompts. This is hard for most existing methods as they only support producing pre-specified head or half-body motion aligned with audio inputs. In this paper, we propose a versatile human animation method, i.e., VersaAnimator, which generates whole-body talking human from arbitrary portrait images, not only driven by audio signal but also flexibly controlled by text prompts. Specifically, we design a text-controlled, audio-driven motion generator that produces whole-body motion representations in 3D synchronized with audio inputs while following textual motion descriptions. To promote natural smooth motion, we propose a code-pose translation module to link VAE codebooks with 2D DWposes extracted from template videos. Moreover, we introduce a multi-modal video diffusion that generates photorealistic human animation from a reference image according to both audio inputs and whole-body motion representations. Extensive experiments show that VersaAnimator outperforms existing methods in visual quality, identity preservation, and audio-lip synchronization.
Machine Learning, ICML
1 Introduction
Talking human video generation aims to animate a portrait (or an avatar) from control signals such as audio, text and pose sequences, producing realistic and natural talking human. This has emerged as a prominent and rapidly growing research direction within video generation, with a wide range of applications in education (Seeger et al., 2021), healthcare (Seitz et al., 2022) and live streaming (Li et al., 2021a), etc..
@@ 1 Introduction
Figure 1: Given a reference image in the first column and an audio clip, our method generates photorealistic talking videos of the person. As demonstrated in the synthesized images, our approach supports arbitrary reference images, i.e., semi-body in the upper demo and whole-body in the lower, and allows users to control or edit the body motion by text prompts, e.g. waving different hands at the closing.
Existing works start from synthesizing talking head for audio-driven speaker video generation (Tian et al., 2025; Wang et al., 2024a; Xu et al., 2024b), where the focuses of human motion generation are primarily on the lip and facial expressions within the head-shoulder region. To enhance the expressiveness of a speaker, later efforts (Corona et al., ; Lin et al., 2024; Meng et al., 2024) extend the animation generation to include the upper body, enabling natural and gestural movements.
These methods can well handle some formal scenarios, such as news broadcasting or speeches, where the generated body motions are typically restricted to the gestures of a reference figure, only involving the upper body without the lower body. Actually, this is quite restrictive, as in talk shows or crosstalk performances, whole-body human motion is essential to impress audiences with infectious speech. Besides, since persons seldom hold still during speaking, whole-body motion shall be generated even for a half-body reference image. Nevertheless, how can we generate whole-body motion for an arbitrary reference 1/4, 1/2 headshot or a whole body portrait, truly extending a talking head to talking human?
The whole-body motion of an audio-driven talking human needs to synchronize with the given audio control signal. Though audio signal can precisely determine lip motion, it is hard to associate audio with specific body motion, since there are multiple ways to convey the same message by body language in different cultures or by different persons. Therefore, this inherent diversity requires users to specify body motion in a speech. In prior work, users have no control over the character’s action, e.g., requiring him to wave goodbye with the left hand or right hand to audiences around the closing words in Figure 1. Thus, how can we allow users to use convenient text prompts to control or edit the body motion generation flexibly?
In this paper, we would like to generate a whole-body talking human from a single reference figure image, either a headshot or portrait, and allow users to employ multimodality controls, i.e., audio signal plus text prompts, to flexibly control or edit the body motion generation. Therefore, we propose a versatile audio-driven talking human generation method, abbreviated as VersaAnimator. This animator consists of two key components: a motion generator that produces pose sequences synchronized with the given audio signal while following the text prompts of body motion; and a multimodal diffusion module that generates talking human video based on the reference image, no matter a 1/4, 1/2 or full-body portrait. We argue that if the input reference only involves upper body, the generation of the whole-body motion still contributes to natural head or upper-body movements during speaking, even only the upper-body regions are rendered as required by users.
Specifically, VersaAnimator generates human pose sequences in 3D first and maps them back to 2D pose sequences later, since 3D human motion offers precise control over the movement of bones and joins. Besides, 3D motion datasets (Guo et al., 2022a) are readily accessible to learn the text-to-motion semantic association. Afterwards, we construct a set of motion tokens to unify the motion representation across different datasets by training a VQ-AVE. To support the multimodal control, we design a two-branch transformer that generates motions conditioned on both audio and text prompts. We further design a new code-pose translation module to synthesize natural and smooth motion when mapping 3D motions to 2D pose sequences, by constructing a relation bank that links the VQ-VAE codebook to 2D poses. Finally, we employ a multimodal diffusion model that incorporate both audio control signal and the 2D pose sequence to produce a talking human aligned with the audio lip sync and expressions. This diffusion model is trained in two stages, learning the whole-body movement first and then the audio-driven facial movement generation.
@@
We have constructed a 40-hour human animation training set that spans from head to whole-body human videos, collected from the internet. Extensive qualitative and quantitative experiments demonstrate that our method achieves state-of-the-art performance in terms of both robustness and naturalness across a wide range of reference images. In addition, our VersaAnimator exhibits excellent text-controlled motion generation capability on the HumanML3D benchmark (Guo et al., 2022a). To our best knowledge, this is the first work dedicated to generating whole-body talking human, with the following contributions:
  • We propose to generate whole-body motions for arbitrary refence images, with no restriction on the visible body region in the reference. The generation of lower body movements enhances the expressiveness of a talking human, even only half-body videos are rendered if required by users, by presenting subtle sway in sync with natural talking rhythm.
  • We design a text-controlled, audio-driven motion generator that produces pose sequences synchronized with the given audio signal, while allowing users to flexibly control or edit body motion generation using convenient text prompts.
  • We design a multimodal diffusion model with two-stage training process for human video generation, learning the whole-body movement first and then the audio-driven facial movement generation.
@@
Figure 2: Overview of our VersaAnimator. The Motion Genration process uses both audio and text modalities to specify the facial and body motion to generate. Then Condition Synergy process generates the pose condition and audio condition as the control signal in video generation. Finally, we inject both conditions into the diffusion model to animate the reference character for Video Generation.
2 Related work
@@
Human Video Generation. In terms of the target region, existing works have focused on either head or body. Remarkable efforts (Tian et al., 2025; Wang et al., 2024a; Xu et al., 2024b; Chen et al., 2024; Zhang et al., 2023c; Ma et al., 2023; Wang et al., 2021; Yin et al., 2022; Prajwal et al., 2020) have concentrated on audio-driven speaker video generation, primarily focusing on the head-shoulder region especially facial expressions, in speech-driven scenarios. Recent efforts (Zhang et al., 2024b; Zhu et al., 2025; Wang et al., 2024b; Feng et al., 2023; Xu et al., 2024c; Hu, 2024; Chang et al., 2023; Karras et al., 2023; Ma et al., 2024) have focused on driving character animation through pose guidance. To enhance expressiveness and realism, VLOGGER (Corona et al., ) and CyberHost (Lin et al., 2024) generate half-body talking videos with movements using only audio and reference maps as inputs. EchoMimicV2 (Meng et al., 2024) need for gesture movements as input in addition to audio. Despite recent advancements, they still exhibit several limitations, such as quite limited body motion and a lack of fine-grained control of gesture movements in the generated videos.
@@
Human Motion Generation. Human motion generation can be broadly categorized into two primary approaches w.r.t. the control inputs: 1) motion synthesis without conditions (Raab et al., 2022; Tevet et al., 2022; Zhang et al., 2020) and 2) motion synthesis with specified multimodal conditions, such as action labels (Xu et al., 2022; Lee et al., 2022; Dou et al., 2023), textual descriptions (Wang et al., ; Chen et al., 2022; Lu et al., 2023), or audio and music (Li et al., b; Tseng et al., 2022; Li et al., a; Dabral et al., 2022). Due to its user-friendly nature and the convenience of language input, text-to-motion is one of the most important motion generation tasks. Given the remarkable success of diffusion-based generative models on AIGC tasks (Rombach et al., 2022), some approaches have employed conditional diffusion models for human motion generation (Zhang et al., 2024a; Chen et al., 2023). Other works (Zhang et al., 2023a; Guo et al., 2024; Wang, 2023) first discretize motions into tokens using vector quantization (Van Den Oord et al., 2017) and then predict the code sequence of motion.
3 Method
3.1 Preliminary and Overview
@@
3D human motions are highly controllable with a rich dataset available (Guo et al., 2022a), containing a variety of motion data paired with corresponding textual descriptions. This motivates us to utilize 3D motions as the whole-body motion representation. On one hand, this enhances the authenticity and plausibility of human motions in videos. On the other hand, by leveraging the rich text-3D motion correspondence data, we can enable users to use convenient text prompts to control or edit the body motion generation flexibly in generated video.
@@
3.2 Audio-driven Motion Generation with Text Control
@@
3.2.1 3D Human Motion Tokenizer
Vector Quantized Variational Autoencoders (VQ-VAE) (Van Den Oord et al., 2017) enable the learning of discrete representations, offering significant advantages for content compression and generation. VQ-VAE reconstructs the motion sequence using an autoencoder with a learnable $K$-size codebook $C=\{c_{k}\}_{k=1}^{K}$, where each code $c_{k}\in\mathbb{R}^{d_{c}}$ corresponds to a discretized feature, and $d_{c}$ denotes the dimension of the discretized feature. Given a motion sequence $M=\{m_{t}\}_{t=1}^{T}$, the encoder $\mathcal{E}$ maps it into a latent feature sequence $Z=\{z_{i}\}_{i=1}^{T/l}$, where $z_{i}\in\mathbb{R}^{d_{c}}$ and $l$ is the temporal downsampling rate of $\mathcal{E}$. For each latent feature $z_{i}$, quantization is performed by selecting the closest element in $C$, resulting in the quantized feature sequence $\hat{Z}=\{\hat{z}_{i}\}_{i=1}^{T/l}$, as follows:
@@
3.2.2 Motion Generation Architecture
@@
We implement the transformer $p_{\theta}^{audio}(z|audio)$ as the primary audio-to-motion architecture. The goal is to predict the code sequence (responded to motion tokens in codebook) conditioned by the audio signal. Specifically, the audio tokens are fed into the transformer encoder to capture long-range dependencies and contextual relationships within the audio sequence. The outputs of each transformer layer are gathered and denoted by $\{f_{s}^{audio}\}_{s=1}^{S}$, where $f^{audio}_{s}$ represent the $s$-th layer’s output and $S$ is set to 8. Then utilizing simple linear layers, we convert last feature $f_{S}^{audio}$ to code probability $I^{\text{audio}}\in\mathbb{R}^{T\times K}$, where $T$ and $K$ represent the predicted motion length and size of the codebook respectively. At timestamp $t$, the probability distribution of the motion codes is as follows:
$P(m_{t}=c_{k})=p_{k}^{t},\ \ k=1,...,K.$ (1)
Then, we select the code sequence $\{c_{i}\}_{i=1}^{T/l}$ that matches the audio signal through the probability distribution $I^{audio}$. Finally, a well-trained decoder and corresponding codebook convert the code sequence into the motion ${M}=\{m_{t}\}_{t=1}^{T}$.
Multimodel Conditions To edit the motion of animated character with text, i.e., to allow the text to control the prediction of the motion token, we add an extra text control branch $p_{\theta}^{text}(z|text)$. Specifically, we learn a text-to-motion transformer to model the tokens conditioned on text. The text-to-motion transformer has the same architecture as the primary audio-to-motion model. Given the text signal, we use CLIP (Radford et al., 2021) for extracting text features and feed into text-to-motion transformer and obtain $\{f_{s}^{text}\}_{s=1}^{S}$, where $f^{text}_{s}$ represents the $s$-th layer’s output. We then consolidate $p_{\theta}^{audio}(z|audio)$ and $p_{\theta}^{text}(z|text)$ by summing them layer by layer.
Figure 3: Structure of the motion generator. Left: The primary audio-to-motion architecture. Right: The two-branch transformer that generates motions conditioned on both audio and text prompts.
3.2.3 Training
@@
$\mathcal{L}_{mdr}=\|M-\hat{M}\|_{1}+\beta\left\|Z-\operatorname{sg}\left[\hat{Z}\right]\right\|_{2}^{2},$ (2)
where $\operatorname{sg}[\cdot]$ denotes the stop-gradient operation, and $\beta$ a weighting factor for embedding constraint. The codebooks are updated via exponential moving average and codebook reset following (Zhang et al., 2023a).
Training Motion Generation Model. We first train the text to motion branch, then freeze the weights and train the audio to motion branch. For text to motion branch, we randomly mask out the sequence elements, by replacing the tokens with a special mask token. The goal is to predict the masked tokens given text. We use CLIP (Radford et al., 2021) for extracting text features. Our training goal is to predict the masked tockens. We directly maximize the log-likelihood of the data distribution $p(\hat{M}\mid text)$:
$\mathcal{L}_{\text{tran }}=\mathbb{E}_{\hat{M}\sim p(\hat{M})}[-\log p(\hat{M}\mid text)],$ (3)
@@
$p(\hat{M}\mid text)=\prod_{i=1}^{T/l}\left(p\left(\hat{m}_{t}\mid text\right)\cdot\left(1-[mask]_{t}\right)+[mask]_{t}\right),$ (4)
@@
$\mathcal{L}_{\text{res }}=\sum_{i=1}^{n}-\log p_{\phi}\left(m_{i}\mid a_{i},text\right),$ (5)
where $m_{i}$ and $a_{i}$ are the $i$-th motion and audio token respectively.
3.2.4 Code-pose Translation Construction
After obtaining the code sequence (motion tokens), the next step is to map them to 2D pose sequences for video generation. Directly mapping 3D human model to 2D pose often results in stiff and unrealistic motions. To address this issue, we construct a relation bank that links the VQ-VAE codebook to 2D poses. Specifically, we extract the SMPL-X data (Cai et al., 2024) and 2D dwpose (Yang et al., 2023) of the template videos, and use the previously trained 3D human motion tokenizer to convert them into code sequences. This allows us to establish the correspondence between the code and 2d pose. By extracting real-world poses from template videos and aligning them with the codes in the codebook, we can create a large set of code-2D pose pairs, thereby enhancing the realism of the generated motions.
3.3 Generating Photorealistic Talking and Moving Humans with Audio and Motion Control

### v3

# Versatile Multimodal Controls for Expressive Talking Human Animation

@@
Abstract.
In filmmaking, directors typically allow actors to perform freely based on the script before providing specific guidance on how to present key actions. AI-generated content faces similar requirements, where users not only need automatic generation of lip synchronization and basic gestures from audio input but also desire semantically accurate and expressive body movement that can be “directly guided” through text descriptions. Therefore, we present VersaAnimator, a versatile framework that synthesizes expressive talking human videos from arbitrary portrait images. Specifically, we design a motion generator that produces basic rhythmic movements from audio input and supports text-prompt control for specific actions. The generated whole-body 3D motion tokens can animate portraits of various scales, producing talking heads, half-body gestures and even leg movements for whole-body images. Besides, we introduce a multi-modal controlled video diffusion that generates photorealistic videos, where speech signals govern lip synchronization, facial expressions, and head motions while body movements are guided by the 2D poses. Furthermore, we introduce a token2pose translator to smoothly map 3D motion tokens to 2D pose sequences. This design mitigates the stiffness resulting from direct 3D to 2D conversion and enhances the details of the generated body movements. Extensive experiments shows that VersaAnimator synthesizes lip-synced and identity-preserving videos while generating expressive and semantically meaningful whole-body motions.
Human animation, Multimodel video generation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; 2025; isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologiesccs: Computing methodologies Artificial intelligence
@@ Abstract.
Figure 1. Given a reference image in the first column and an audio clip, our method generates photorealistic talking videos of the person. As demonstrated in the synthesized images, our approach supports arbitrary reference images, i.e., semi-body in the upper demo and whole-body in the lower, and allows users to control or edit the body motion by text prompts, e.g. waving different hands at the closing.
[…]
Enjoying the baseball game from the third-base seats. Ichiro Suzuki preparing to bat.
1. Introduction
Recent advances in diffusion models have significantly inspired research in the domains of talking head (Tian et al., 2025; Wang et al., 2024b; Xu et al., 2024a) and human animation (Zhang et al., 2024a; Zhu et al., 2025; Wang et al., 2024a), enhancing the quality and expressiveness of one-shot video generation. Recently, some innovative works (Corona et al., [n. d.]; Lin et al., 2024; Meng et al., 2024; Lin et al., 2025) successfully attempt to integrate these two directions, achieving synchronized generation of facial expressions and body movements, evolving talking heads into talking humans.
However, even state-of-the-art methods only demonstrate the potential to support some AI-generated content (AIGC) scenarios through demo videos. The quality of their generated videos and product functionality still struggle to meet the demands of commercial-grade scenarios. The limitations of existing methods primarily manifest in three aspects:
(1) Gestures generated from audio are simple, lacking expressiveness and semantic correspondence. Some methods generate random speech gestures based solely on audio, resulting in repetitive gestures with limited motion range that fail to convey sufficient semantic information.
(2) Generated videos are not editable, including facial expressions and body movements. This leaves users with no direct means to adjust the results apart from regeneration, making outcomes entirely random. Achieving desired results becomes challenging and potentially time-consuming.
(3) The supported portrait scale is limited. Current methods mostly focus on gesture generation for half-body reference images, lacking the ability to generate corresponding whole-body movements, especially leg movements, for whole-body photographs. This limitation significantly constrains the range of producible video content.
Addressing these issues, we delve into their underlying causes and explore potential solutions. Unlike the direct correspondence between speech and lip movements, the relationship between body movements and speech content is a fuzzy many-to-many mapping. Certain specific, large-amplitude gestures are not entirely determined by speech alone but are also related to personal habits and contextual content. In film production, directors typically allow actors to perform freely based on the script before providing specific guidance on how to present key actions. AIGC faces similar requirements, where users not only need automatic generation of lip synchronization and basic gestures from audio input but also desire semantically accurate body movements that can be “directly guided” through text descriptions. Therefore, we argue that a reasonable motion generation model should use audio input to provide basic rhythmic movements, while expressive and semantically explicit actions should support user control through means such as text descriptions. Additionally, we believe that movements should be generated based on whole-body motion representations. The range of motion generation should not be limited by the reference image, instead, the generated motions should be adaptable to and capable of driving reference images of any scale.
Based on these considerations, we propose VersaAnimator, a versatile human animation framework that generates expressive talking human videos from arbitrary portrait images, not only driven by audio signals but also flexibly controlled by text prompts. Specifically, we design a motion generator that produces basic rhythmic movements from the audio input and supports text-prompt control for specific actions. The generated whole-body 3D motion tokens can animate portraits of various scales, producing talking heads, half-body gestures and even leg movements for whole-body images. We utilize large-scale 3D motion datasets (Guo et al., 2022a) to facilitate the learning of semantic associations between text descriptions and motion. A Vector Quantized Variational Autoencoder (VQ-VAE) is trained to unify motion representations across diverse datasets to a set of motion tokens. To enable multimodal control, we design a dual-branch transformer architecture that generates motion tokens conditioned on both audio signals and text prompts. Besides, we introduce a multi-modal controlled video diffusion that generates photorealistic videos, where speech signals govern lip synchronization, facial expressions, and head motions while body movements are guided by the 2D poses. This diffusion model is trained in two stages, first learning the pose-to-video capability at different scales, and then the audio-driven facial motion generation. We also design a token-to-pose translator to smoothly map 3D motion tokens to 2D pose sequences. This design mitigates the stiffness resulting from direct 3D to 2D conversion and enhances the details of the generated body movements.
@@ 1. Introduction
We have constructed a 40-hour human animation training set that spans from head to whole-body human videos. Extensive qualitative and quantitative experiments demonstrate that our method achieves state-of-the-art performance, synthesizing lip-synced and identity-preserving videos while generating expressive and semantically meaningful whole-body motions. Our work has the following main contributions:
  • We propose a versatile talking human framework which, to our knowledge, is the first to simultaneously achieve audio-driven, text-controlled and whole-body scale motion generation, offering diverse application scenarios.
  • We propose a motion generation approach where audio controls basic motion rhythm and text controls semantically rich actions, and implement the generation of whole-body level 3D motion tokens conditioned on both audio signals and text prompts, capable of animating portrait images of various scales.
  • We design a multimodal video diffusion that can simultaneously control the fine-grained generation of facial expressions and body movements through audio and pose sequences, achieving high-quality and expressive human video generation.
@@
Figure 2. Overview of our VersaAnimator. The Motion Generation process uses both audio and text modalities to specify the facial and body motion to generate. Then Condition Synergy process generates the pose condition and audio condition as the control signal in video generation. Finally, we inject both conditions into the diffusion model to animate the reference character for Video Generation.
[…]
2. Related work
@@
Human Video Generation. In terms of the target region, existing works have focused on either head or body. Remarkable efforts (Tian et al., 2025; Wang et al., 2024b; Xu et al., 2024a; Chen et al., 2024; Zhang et al., 2023a; Ma et al., 2023; Wang et al., 2021; Yin et al., 2022; Prajwal et al., 2020) have concentrated on audio-driven speaker video generation, primarily focusing on the head-shoulder region especially facial expressions, in speech-driven scenarios. Recent efforts (Zhang et al., 2024a; Zhu et al., 2025; Wang et al., 2024a; Feng et al., 2023; Xu et al., 2024c; Hu, 2024; Chang et al., 2023; Karras et al., 2023; Ma et al., 2024) have focused on driving character animation through pose guidance. To enhance expressiveness and realism, VLOGGER (Corona et al., [n. d.]) and CyberHost (Lin et al., 2024) generate half-body talking videos with movements using only audio and reference maps as inputs. EchoMimicV2 (Meng et al., 2024) need for gesture movements as input in addition to audio. Despite recent advancements, they still exhibit several limitations, such as quite limited body motion and a lack of fine-grained control of gesture movements in the generated videos.
@@
Human Motion Generation. Human motion generation can be broadly categorized into two primary approaches w.r.t. the control inputs: 1) motion synthesis without conditions (Raab et al., 2023; Tevet et al., 2022; Zhang et al., 2020) and 2) motion synthesis with specified multimodal conditions, such as action labels (Xu et al., 2023; Lee et al., 2023; Dou et al., 2023), textual descriptions (Wang et al., 2022; Chen et al., 2023; Lu et al., 2023), or audio and music (Li et al., 2024a; Tseng et al., 2023; Li et al., 2022; Dabral et al., 2023). Due to its user-friendly nature and the convenience of language input, text-to-motion is one of the most important motion generation tasks. Given the remarkable success of diffusion-based generative models on AIGC tasks (Rombach et al., 2022), some approaches have employed conditional diffusion models for human motion generation (Zhang et al., 2024b; Chen et al., 2023). Other works (Zhang et al., 2023c; Guo et al., 2024; Wang, 2023) first discretize motions into tokens using vector quantization (Van Den Oord et al., 2017) and then predict the code sequence of motion.
3. Method
3.1. Preliminary and Overview
@@
3D human motions are highly controllable with a rich dataset available (Guo et al., 2022a), containing a variety of motion data paired with corresponding textual descriptions. This motivates us to utilize 3D motions as the whole-body motion representation. On one hand, this enhances the authenticity and plausibility of human motions in videos. On the other hand, by leveraging the rich text-3D motion correspondence data, we can enable users to use convenient text prompts to control or edit the body motion generation flexibly in the generated video.
@@
3.2. Audio-driven Motion Generation with Text Control
@@
3.2.1. 3D Human Motion Tokenizer
Vector Quantized Variational Autoencoders (VQ-VAE) (Van Den Oord et al., 2017) enable the learning of discrete representations, offering significant advantages for content compression and generation. VQ-VAE reconstructs the motion sequence using an autoencoder with a $K$-size learnable codebook $C=\{c_{k}\}_{k=1}^{K}$, where $c_{k}\in\mathbb{R}^{d_{c}}$ represents the discretized motion token and $d_{c}$ denotes the feature dimension. Given a motion sequence $M=\{m_{t}\}_{t=1}^{T}$, the encoder $\mathcal{E}$ maps it into a latent feature sequence $Z=\{z_{i}\}_{i=1}^{T/l}$, where $z_{i}\in\mathbb{R}^{d_{c}}$ and $l$ is the temporal downsampling rate of $\mathcal{E}$. For each latent feature $z_{i}$, quantization is performed by selecting the closest element in $C$, resulting in the quantized feature sequence $\hat{Z}=\{\hat{z}_{i}\}_{i=1}^{T/l}$, as follows:
@@
3.2.2. Motion Generation Architecture
@@
We implement the transformer $p_{\theta}^{audio}(\hat{z}|audio)$ as the primary audio-to-motion architecture, where $\hat{z}$ denotes the predicted motion token here. The goal is to predict the code sequence (responded to motion tokens in codebook) conditioned by the audio signal. Specifically, the audio tokens are fed into the transformer encoder to capture long-range dependencies and contextual relationships within the audio sequence. The outputs of each transformer layer are gathered and denoted by $\{f_{s}^{audio}\}_{s=1}^{S}$, where $f^{audio}_{s}$ represent the $s$-th layer’s output and $S$ is set to 8. Then utilizing simple linear layers, we convert last feature $f_{S}^{audio}$ to code probability $I^{\text{audio}}\in\mathbb{R}^{T\times K}$, where $T$ and $K$ represent the predicted motion length and size of the codebook respectively. At timestamp $t$, the probability distribution of the motion tokens in codebook is as follows:
(1) $P(c_{k})=p_{k}^{t},\ \ k=1,...,K.$
We then select the code sequence that aligns with the audio signal based on the probability distribution $I^{audio}$, and use the codebook to retrieve the corresponding motion tokens $\{\hat{z}_{i}\}_{i=1}^{T/l}$. Finally, a trained decoder converts them into the motion sequence ${\hat{M}}=\{\hat{m}_{t}\}_{t=1}^{T}$.
Multimodel Conditions. To edit the motion of animated character with text, i.e., to allow the text to control the prediction of the motion token, we adopted a two-branch transformer $p_{\theta}(\hat{z}|text,audio)$. Specifically, we learn a text-to-motion transformer to model the tokens conditioned on text. The text-to-motion transformer has the same architecture as the primary audio-to-motion model. Given the text signal, we use CLIP (Radford et al., 2021) for extracting text features and feed into text-to-motion transformer and obtain $\{f_{s}^{text}\}_{s=1}^{S}$, where $f^{text}_{s}$ represents the $s$-th layer’s output. We then consolidate two transformer structure by summing them layer by layer as show in Figure 3 .
Figure 3. Structure of the motion generator. Left: The primary audio-to-motion architecture. Right: The two-branch transformer that generates motions conditioned on both audio and text prompts.
[…]
3.2.3. Training
@@
(2) $\mathcal{L}_{mdr}=\|M-\hat{M}\|_{1}+\beta\left\|Z-\operatorname{sg}\left[\hat{Z}\right]\right\|_{2}^{2},$
where $\operatorname{sg}[\cdot]$ denotes the stop-gradient operation, and $\beta$ a weighting factor for embedding constraint. The codebooks are updated via exponential moving average and codebook reset following (Zhang et al., 2023c).
Training Motion Generation Model. We first train the text to motion branch, then freeze the weights and train the audio to motion branch. For text to motion branch, we randomly mask out the sequence elements, by replacing the tokens with a special mask token. The goal is to predict the masked tokens given text. We use CLIP (Radford et al., 2021) for extracting text features. Our training goal is to predict the masked tockens. We directly maximize the log-likelihood of the data distribution $p(\hat{Z}\mid text)$:
(3) $\mathcal{L}_{\text{tran }}=\mathbb{E}_{\hat{Z}\sim p(\hat{Z})}[-\log p(\hat{Z}\mid text)],$
@@
(4) $p(\hat{Z}\mid text)=\prod_{i=1}^{T/l}\left(p\left(\hat{z}_{t}\mid text\right)\cdot\left(1-[mask]_{t}\right)+[mask]_{t}\right),$
@@
(5) $\mathcal{L}_{\text{audo }}=\sum_{i=1}^{n}-\log p_{\phi}\left(\hat{z}_{i}\mid a_{i},text\right),$
where $\hat{z}_{i}$ and $a_{i}$ are the $i$-th motion and audio token respectively.
3.2.4. Token2pose Translator Construction
After obtaining the generated motion tokens, the next step is to map them to 2D pose sequences for video generation. Directly mapping 3D human model to 2D pose often results in stiff and unrealistic motions. To address this issue, we construct a relation bank that links all motion tokens to detailed 2D poses. Specifically, we collect whole-body template videos covering various types of actions, simultaneously extracting normalized 2D pose and SMPL-X data. Next, we use the previously trained 3D human motion tokenizer to convert SMPL-X data into motion tokens. By extracting real-world poses from template videos and aligning them with motion tokens, we can create a large set of token2pose pairs. We use these pairs to map the generated motion tokens to real 2D poses and further adapt them to the target identity, thereby enhancing the realism of the generated motions.
3.3. Generating Photorealistic Talking and Moving Humans with Audio and Motion Control